OpenStudy (anonymous):

need help ? 8x - 1 = -4x + 23 8x - 1 = -4x + 92

hero (hero):

Hint: 8x - 1 = 8x - 1

hero (hero):

Oh this is an interesting one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

i thought i would have to distribute

OpenStudy (skullpatrol):

As Hero pointed out 8x - 1 = 8x - 1 right?

hero (hero):

That's not the answer. But both -4x + 23 and -4x + 92 are both equal to the same number.

hero (hero):

So you have to set them equal then solve for x (if possible): -4x + 23 = -4x + 92

OpenStudy (skullpatrol):

@Markitos Are you still there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im just confused srry

OpenStudy (anonymous):

so solve for -4x + 23 = -4x + 92

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Please ask questions right away if you don't understand, okay?

OpenStudy (anonymous):

so do i solve for -4x + 23 = -4x + 92